# MAGPIE: Multilevel–Adaptive–Guided solver for ptychographic phase retrieval

### v4

@@ 2 Surrogate Minimization
In this section, we introduce a surrogate objective for Eqn. (1). This surrogate not only majorizes Eqn. (1) (see Section 2.1 for a precise definition of majorization) and iteratively recovers its optimal solution, but is also sufficiently general to encompass the entire PIE family of solvers. An equivalent formulation, presented in [23], has demonstrated superior performance on the general phase-retrieval problem.
@@
$\widetilde{\Phi}\left(\bm{z};\bm{z}^{(j)}\right)$ $=\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}P_{k}\bm{z}-\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)\right\|_{2}^{2}\,,$ (7)
$=\sum_{k=1}^{N}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)=\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right\|_{2}^{2}\,.$
where $\mathcal{R}_{k}$ is defined in Definition 1.1. We note that in the definition of $\mathcal{R}_{k}$, the term $\left[\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right)\right]_{r}$ is ill-defined if $\left[\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right]_{r}=0$ at some index $r$, so that we complement the original definition with the following specification:
$\left[\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right)\right]_{r}=\begin{cases}0\,,&\text{if }j=0\,\&\left[\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right]_{r}=0,\\ \left[\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j-1)}\right)\right)\right]_{r}\,,&\text{if }j\geq 1\,\&\left[\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right]_{r}=0,\\ \left[\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right)\right]_{r}\,,&else\end{cases}$ (8)
@@
$\bm{z}_{k}^{(j)}\leftarrow\bm{z}_{k}^{+}=\operatornamewithlimits{argmin}_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)+\frac{1}{2}\bm{u}\cdot\left|\bm{z}_{k}-\bm{z}_{k}^{(j)}\right|^{2}\,.$ (11)
Notice that the update formulation  Eqn. (11) is equivalent to applying one step of the proximal point method [24] to  Eqn. (10).
@@ 2.1 Properties of the surrogate model
Our proposed surrogate model in Eqn. (9) enjoys many desirable properties and is designed in accordance with the general principle of majorization [25] (see Figure 2).
Figure 2: Illustration of the majorization property adapted from [26].
@@
The proof of this theorem appears in Appendix B.1. The first statement is straightforward to verify. To show the inequality (12), we deploy the definitions in Eqn. (1) and Eqn. (7) and prove the inequality for every $\Phi_{k}$. The last statement comes from straightforward computation. We call $\widetilde{\Phi}$ a majorization of $\Phi$ due to the inequality Eqn. (12).
Furthermore, minimizing the surrogate (see  Eqn. (9)) provides a descent direction for the original objective function, as implied in the following theorem:
Theorem 2.2 (Convergence and optimality).
@@
We leave the proof of this theorem to Appendix B.2. The monotonicity inequality follows directly from Proposition 2.1, and the first-order optimality of $\Phi$ follows from the corresponding first-order optimality condition on the quadratic surrogate, with $\mathcal{R}_{k}$ specified in (8)
The convergence rate of the surrogate minimization using  Eqn. (9) is analyzed in the following theorem:
Theorem 2.3 (Convergence rate).
@@
$\min_{0\leq j\leq n}\left\|\nabla_{\bm{z}}\Phi\left(\bm{z}^{(j)}\right)\right\|_{2}^{2}\leq\frac{\left\|\sum_{k=1}^{N}P_{k}^{\top}|\bm{Q}|^{2}\right\|_{\infty}}{2(n+1)}\Phi\left(\bm{z}^{(0)}\right)\,,$
where $\sum_{k=1}^{N}P_{k}^{\top}|\bm{Q}|^{2}\in\mathbb{C}^{n^{2}}$ and $\left\|\sum_{k=1}^{N}P_{k}^{\top}|\bm{Q}|^{2}\right\|_{\infty}$ is the maximum magnitude among its entries.
We leave the proof of this theorem to Appendix C. The overall derivation follows that in [27], where the increments at each iteration are bounded, and a telescoping-sum argument is applied.
@@ 3 MAGPIE
As demonstrated in [20], the multigrid solver improves computational efficiency, while the surrogate majorization model provides favorable convergence properties. We now integrate these approaches into a single, tractable, and robust method, which we term the Multilevel–Adaptive–Guided Ptychographic Iterative Engine (MAGPIE). The name reflects its multigrid structure, adaptive construction of coarse-grid surrogates, and guided fine-grid reconstruction via coarse-grid search directions. We present the algorithm in Section 3.1 and defer discussion of its well-definedness, consistency, descent properties, and automatic regularization selection to Section 3.2.
3.1 Solver development of MAGPIE
@@
$\min_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)=\frac{1}{2}\left\|\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right\|_{2}^{2}\,,$ (13)
@@
$\min_{\bm{z}_{H,k}}\widetilde{\Phi}_{H,k}\left(\bm{z}_{H,k};\bm{z}_{k}^{(j)}\right)=\frac{1}{2}\left\|\bm{Q}_{H}\bm{z}_{H,k}-\mathcal{R}_{H,k}\left(\bm{z}_{k}^{(j)}\right)\right\|_{2}^{2}\,.$ (14)
@@
  • $\bm{Q}_{H}=\bm{I}_{h}^{H}\bm{Q}\in\mathbb{C}^{(m/2)^{2}}$: downsampled probe by directly applying the restriction operator.
@@
Both the fine- and coarse-grid subproblems are quadratic and, in principle, admit closed-form solutions. In practice, however, a direct solve is ill-posed: whenever entries of $\bm{Q}$ or $\bm{Q}_{H}$ vanish, the corresponding normal equations lose rank and the solution is no longer unique. As with the full surrogate problem in Eqn. (10), we therefore replace an exact solve by a single proximal-point (relaxation) step with fine-grid regularization term $\bm{u}$ and coarse-grid regularization $\bm{u}_{H}$. Concretely, we define:
  • $\bm{W}_{\bm{z}}\in\mathbb{R}^{m^{2}}_{\geq 0}$: constant weight vector for downsampling the current object estimate $\bm{z}_{k}^{(j)}$.
  • $\bm{z}_{H,k}^{(j)}=\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\odot\bm{z}_{k}^{(j)}\right)$: downsampled current object estimate by weighted average pooling.
  • $\bm{W}_{\bm{u}}^{H}\in\mathbb{R}^{(m/2)^{2}}_{\geq 0}$: constant weight vector for downsampling regularization term $\bm{u}$.
@@
$\hat{\bm{z}}_{H,k}^{(j)}=\bm{z}_{H,k}^{(j)}+\frac{\overline{\bm{Q}_{H}}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\left(\mathcal{R}_{H,k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}_{H}\bm{z}_{H,k}^{(j)}\right)\,,$
@@
$\bm{z}_{k}^{(j)}\leftarrow\bm{z}_{k}^{+}=\widetilde{\bm{z}}_{k}^{(j)}+\frac{\overline{\bm{Q}}}{\bm{u}+|\bm{Q}|^{2}}\left(\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}\widetilde{\bm{z}}_{k}^{(j)}\right)\,.$
For computational purposes, the weights ($\bm{W}_{\bm{z}}$, $\bm{W}_{\mathcal{R}}$, and $\bm{W}_{\bm{u}}$) translate information across levels and must be defined carefully to ensure compatibility with the fine-grid optimization. We present the weight formula in Eqn. (16). These definitions guarantee well-definedness, consistency, descent directions, and stability. We discuss these properties in detail in Section 3.2.
We term this multigrid solver for each surrogate model the Multilevel–Adaptive–Guided Proximal Solver (MAGPS). Specifically, “Adaptive” denotes the automatic construction of coarse-grid surrogate terms using the weights specified in Eqn. (16) without manual tuning, “Guided” emphasizes the directional coarse-to-fine corrections in Eqn. (15), and “Proximal Solver” highlights the use of proximal-point updates within the multigrid framework. The pseudocode for this solver is summarized in Algorithm 3.
@@
     $\bm{z}^{+}_{k}\leftarrow\bm{z}_{k}^{(j)}+\frac{\overline{\bm{Q}}}{\bm{u}+|\bm{Q}|^{2}}\left(\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}\bm{z}_{k}^{(j)}\right)$.
@@
     Coarse probe: $\bm{Q}_{H}\leftarrow\bm{I}_{h}^{H}\left(\bm{Q}\right)$.
     Set weights for downsampling the object: $\bm{W}_{\bm{z}}\leftarrow\frac{|\bm{Q}|^{2}}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}$.
     Set weights for downsampling the revised exit wave: $\bm{W}_{\mathcal{R}}\leftarrow\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\odot\bm{W}_{\bm{z}}}{\bm{Q}}$.
     Set weights for coarse-grid regularization: $\bm{W}_{\bm{u}}^{H}\leftarrow\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}$.
@@
     $\bm{z}^{+}_{k}\leftarrow\widetilde{\bm{z}}^{(j)}_{k}+\frac{\overline{\bm{Q}}}{\bm{u}+|\bm{Q}|^{2}}\left(\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}\widetilde{\bm{z}}^{(j)}_{k}\right)$.
@@
2:     for $k$ randomly drawn from $1$ to $N$ without replacement do
@@ 3.2 Converging properties of MAGPIE and weight selection
Different choices of the weight vectors lead to different performance of the algorithm. Now we specify our choices of the weight vectors below. These definitions of $\bm{W}_{\bm{z}}$, $\bm{W}_{\mathcal{R}}$, and $\bm{W}_{\bm{u}}^{H}$ ensure the robust performance of MAGPIE.
$\bm{W}_{\bm{z}}=\frac{|\bm{Q}|^{2}}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}\,,\quad\bm{W}_{\mathcal{R}}=\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\odot\bm{W}_{\bm{z}}}{\bm{Q}}=\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\odot|\bm{Q}|^{2}}{\bm{Q}\odot\left(\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}\right)}\,,\quad\bm{W}_{\bm{u}}^{H}=\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\,.$ (16)
As will be shown more clearly in Propositions 3.1, 3.2, and 3.4, and in Theorem 3.3, these specific forms not only prevent singularities, maintain coarse-to-fine consistency, and preserve the descent property, but also ensure stability. It is worth noting that the quantity $\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}$ appears in both $\bm{W}_{\bm{z}}$ and $\bm{W}_{\mathcal{R}}$. Its physical meaning is straightforward: in each $2\times 2$ bin of $|\bm{Q}|^{2}$, $\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}$ computes the average of the four entries and upsamples that average back to the original four positions, whereas $|\bm{Q}|^{2}$ each entry’s original value. As a result,
$\bm{z}_{H,k}^{(j)}=\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\bm{z}_{k}^{(j)}\right)$
@@
To prove this proposition, we realize that $\mathbf{W}_{\mathbf{z}}$ represents the weighted average within a bin and the bin size is $4$. The control for the other two terms comes from Jensen’s inequality used on quadratic functions. The proof is left to Appendix E.
@@
Given the weight vectors defined in  Eqn. (16), the coarse-level objective $\widetilde{\Phi}_{H,k}$ (as defined in  Eqn. (14)) is consistent with the fine-level objective $\widetilde{\Phi}_{k}$ (as defined in  Eqn. (13)) in the sense that, for all $\bm{z}_{k}\in\mathbb{C}^{m^{2}}$, given $\bm{z}_{H,k}=\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\bm{z}_{k}\right)$, we have:
  • $\widetilde{\Phi}_{H,k}\left(\bm{z}_{H,k};\bm{z}_{k}^{(j)}\right)\leq\frac{1}{4}\left\|\bm{W}_{\mathcal{R}}\right\|_{\infty}^{2}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)$
  • $\left\|\nabla_{\bm{z}_{H,k}}\widetilde{\Phi}_{H,k}\left(\bm{z}_{H,k};\bm{z}_{k}^{(j)}\right)\right\|_{2}\leq\frac{1}{2}\left\|\bm{W}_{\bm{u}}^{H}\right\|_{\infty}\,\left\|\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)\right\|_{2}$
We leave the proof to Appendix F. Most of the proof is a straightforward derivation.
Furthermore, the parameters are chosen so that the coarse-grid solution of Eqn. (14), obtained via the proximal point method, provides a descent direction for the fine-level problem. This is formalized in the following theorem.
@@
Given the weight vectors defined in  Eqn. (16), a single proximal-point sweep on the coarse surrogate (14), initialized at $\bm{z}^{\prime}_{H,k}=\bm{I}_{h}^{H}(\bm{W}_{\bm{z}}\bm{z}^{\prime}_{k})$, provides a descent direction for the fine-grid surrogate (13) at $\bm{z}^{\prime}_{k}$. That is:
$\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)^{*}\bm{I}_{H}^{h}\left(\bm{e}_{H,k}\right)\leq 0\,,$
where $\bm{I}_{H}^{h}\left(\bm{e}_{H,k}\right)=\bm{I}_{H}^{h}\left(\widehat{\bm{z}_{H,k}}-\bm{z}_{H,k}^{\prime}\right)$, and $\widehat{\bm{z}_{H,k}}$ solves the proximal point on the coarse level:
$\widehat{\bm{z}_{H,k}}=\operatornamewithlimits{argmin}_{\bm{z}_{H,k}}\widetilde{\Phi}_{H,k}\left(\bm{z}_{H,k};\bm{z}_{k}^{(j)}\right)+\frac{1}{2}\bm{u}_{H}\cdot\left|\bm{z}_{H,k}-\bm{z}_{H,k}^{\prime}\right|^{2}\,.$ (17)

### v6

@@ 3 Surrogate Minimization
In this section, we introduce a surrogate objective for Eqn. (1). This surrogate not only majorizes Eqn. (1) (see Section 3.1 for a precise definition of majorization) and iteratively recovers its optimal solution, but is also sufficiently general to encompass the entire PIE family of solvers. An equivalent formulation, presented in [MM_phaseretrieval], has demonstrated superior performance on the general phase-retrieval problem.
@@
$\widetilde{\Phi}\left(\bm{z};\bm{z}^{(j)}\right)$ $=\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\odot P_{k}\bm{z}-\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)\right\|_{2}^{2}\,$, (7)
$=\sum_{k=1}^{N}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)=\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right\|_{2}^{2}\,.$
where $\mathcal{R}_{k}$ is defined in Definition 2.1. We note that in the definition of $\mathcal{R}_{k}$, the term $\left[\theta\left(\mathcal{F}\left(\bm{Q}\odot\bm{z}_{k}^{(j)}\right)\right)\right]_{r}$ is ill-defined if $\left[\mathcal{F}\left(\bm{Q}\odot\bm{z}_{k}^{(j)}\right)\right]_{r}=0$ at some index $r$, so that we complement the original definition with the following specification [^5] This specification improves the numerical stability of the updates.:
$\left[\theta\left(\mathcal{F}\left(\bm{Q}\odot\bm{z}_{k}^{(j)}\right)\right)\right]_{r}=\begin{cases}0\,,&\text{if }j=0\,\&\left[\mathcal{F}\left(\bm{Q}\odot\bm{z}_{k}^{(j)}\right)\right]_{r}=0,\\ \left[\theta\left(\mathcal{F}\left(\bm{Q}\odot\bm{z}_{k}^{(j-1)}\right)\right)\right]_{r}\,,&\text{if }j\geq 1\,\&\left[\mathcal{F}\left(\bm{Q}\odot\bm{z}_{k}^{(j)}\right)\right]_{r}=0,\\ \left[\theta\left(\mathcal{F}\left(\bm{Q}\odot\bm{z}_{k}^{(j)}\right)\right)\right]_{r}\,,&else\end{cases}$ (8)
@@
$\bm{z}_{k}^{(j)}\leftarrow\bm{z}_{k}^{+}=\operatornamewithlimits{argmin}_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)+\frac{1}{2}\bm{u}^{\top}\left|\bm{z}_{k}-\bm{z}_{k}^{(j)}\right|^{2}\,.$ (11)
Notice that the update formulation  Eqn. (11) is equivalent to applying one step of the proximal point method [Proximal] to  Eqn. (10).
@@ 3.1 Properties of the surrogate model
Our proposed surrogate model in Eqn. (9) enjoys many desirable properties and is designed in accordance with the general principle of majorization [mm_alg_Lange] (see Figure 2).
Figure 2: Illustration of the majorization property adapted from [SMM_graph].
@@
The detailed proof of this proposition appears in Appendix C.1. The first statement is straightforward to verify. To establish inequality (12), we use the definitions in Eqs. (1) and (7) and verify the result for each $\Phi_{k}$. A similar proof can be found in [MM_phaseretrieval]. The proof of the final statement is adapted from [mm_convergence].
Furthermore, minimizing the surrogate (see  Eqn. (9)) provides a descent direction for the original objective function, as implied in the following proposition:
Proposition 3.2 (Convergence and optimality).
@@
The detailed proof of this proposition appears in Appendix C.2. The monotonicity inequality follows directly from Proposition 3.1, and the first-order optimality of $\Phi$ follows from the corresponding first-order optimality condition on the quadratic surrogate, with $\mathcal{R}_{k}$ specified in (8). This is a standard result of a majorization–minimization algorithm [mm_convergence].
The convergence rate of the surrogate minimization using  Eqn. (9) is analyzed in the following proposition:
Proposition 3.3 (Convergence rate).
@@
$\min_{0\leq j\leq n}\left\|\nabla_{\bm{z}}\Phi\left(\bm{z}^{(j)}\right)\right\|_{2}^{2}\leq\frac{\left\|\sum_{k=1}^{N}P_{k}^{\top}|\bm{Q}|^{2}\right\|_{\infty}}{2(n+1)}\Phi\left(\bm{z}^{(0)}\right)\,.$
The detailed proof of this proposition appears in Appendix D. The overall derivation follows that in [mm_convergence], where the increments at each iteration are bounded, and a telescoping-sum argument is applied.
@@ 4 MAGPIE
As demonstrated in [MGOPT_ptycho], the multigrid solver improves computational efficiency, while the surrogate majorization model provides favorable convergence properties. We now integrate these approaches into a single, tractable, and robust method, which we term the Multilevel–Adaptive–Guided Ptychographic Iterative Engine (MAGPIE). The name reflects its multigrid structure, adaptive construction of coarse-grid surrogates, and guided fine‐grid reconstruction via coarse‐grid search directions. We present the algorithm in Section 4.1 and defer discussion of its well-definedness, consistency, descent properties, and automatic regularization selection to Section 4.2.
4.1 Algorithmic Development of MAGPIE
@@
$\min_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)=\frac{1}{2}\left\|\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right\|_{2}^{2}\,$, (13)
@@
$\min_{\bm{z}_{H,k}}\widetilde{\Phi}_{H,k}\left(\bm{z}_{H,k};\bm{z}_{k}^{(j)}\right)=\frac{1}{2}\left\|\bm{Q}_{H}\odot\bm{z}_{H,k}-\mathcal{R}_{H,k}\left(\bm{z}_{k}^{(j)}\right)\right\|_{2}^{2}\,.$ (14)
@@
  • $\bm{Q}_{H}=\bm{I}_{h}^{H}\bm{Q}\in\mathbb{C}^{(m/2)^{2}}$: downsampled probe.
@@
Both the fine- and coarse-grid subproblems are quadratic and, in principle, admit closed-form solutions. In practice, however, a direct solve is ill-posed: whenever entries of $\bm{Q}$ or $\bm{Q}_{H}$ vanish, the corresponding normal equations lose rank, and the solution is no longer unique. As with the full surrogate problem in Eqn. (10), we therefore replace an exact solve with a single proximal-point (relaxation) step that features a fine-grid regularization term $\bm{u}$ and coarse-grid regularization $\bm{u}_{H}$. Concretely, we define:
  • $\bm{W}_{\bm{z}}\in\mathbb{R}^{m^{2}}_{\geq 0}$: constant weight vector for downsampling $\bm{z}_{k}^{(j)}$.
  • $\bm{z}_{H,k}^{(j)}=\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\odot\bm{z}_{k}^{(j)}\right)$: downsampled current object estimate.
  • $\bm{W}_{\bm{u}}^{H}\in\mathbb{R}^{(m/2)^{2}}_{\geq 0}$: constant weight vector for downsampling $\bm{u}$.
@@
$\hat{\bm{z}}_{H,k}^{(j)}=\bm{z}_{H,k}^{(j)}+\overline{\bm{Q}_{H}}\oslash\left(\bm{u}_{H}+|\bm{Q}_{H}|^{2}\right)\odot\left(\mathcal{R}_{H,k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}_{H}\odot\bm{z}_{H,k}^{(j)}\right)\,$,
@@
$\bm{z}_{k}^{(j)}\leftarrow\bm{z}_{k}^{+}=\widetilde{\bm{z}}_{k}^{(j)}+\overline{\bm{Q}}\oslash\left(\bm{u}+|\bm{Q}|^{2}\right)\odot\left(\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}\odot\widetilde{\bm{z}}_{k}^{(j)}\right)\,.$
For computational purposes, the weights ($\bm{W}_{\bm{z}}$, $\bm{W}_{\mathcal{R}}$, and $\bm{W}_{\bm{u}}$) transfer information across levels and must be chosen carefully to ensure compatibility with the fine-grid optimization. We construct these weights in a sequential manner to guarantee consistency, well-definedness, descent directions, and stability.
Consistency
[…]
$\bm{W}_{\mathcal{R}}=(\bm{I}_{H}^{h}\bm{Q}_{H})\odot\bm{W}_{\bm{z}}\oslash\bm{Q}.$
This relation is proved in Proposition 4.2.
Well-definedness
The consistency relation above implies that the expression for $\bm{W}_{\mathcal{R}}$ may suffer from singularities due to division by $\bm{Q}$. To ensure well-definedness, we therefore choose $\bm{W}_{\bm{z}}$ so that these singularities are canceled. There are many possible choices; in this paper, we take
$\bm{W}_{\bm{z}}=|\bm{Q}|^{2}\oslash\left(\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}\right),$
which has a desirable physical interpretation that we discuss in detail in Section 4.2. In addition, Proposition 4.1 shows that all three weights are well defined and introduce no singularities.
Descent direction and stability
We next need to show that the coarse-grid update in Eqn. (15) is well behaved. In Proposition 4.3, we prove that this update yields a descent direction. We then control the magnitude of the update by tuning the regularization term. The regularization must balance stability (which favors a larger value) against the step size (which favors a smaller value). In Proposition 4.4, we show that
$\bm{W}_{\bm{u}}^{H}=|\bm{Q}_{H}|^{2}\oslash\left(\bm{I}_{h}^{H}|\bm{Q}|^{2}\right)$
is a suitable choice.
We term this multigrid solver for each surrogate model the Multilevel–Adaptive–Guided Proximal Solver (MAGPS). Specifically, “Adaptive” denotes the automatic construction of coarse‐grid surrogate terms using the weights specified in Eqn. (16) without manual tuning, “Guided” emphasizes the directional coarse‐to‐fine corrections in Eqn. (15), and “Proximal Solver” highlights the use of proximal‐point updates within the multigrid framework. The pseudocode for this solver is summarized in Algorithm 3.
@@
We have made an implementation of MAGPIE available online [^7] Repository URL: https://github.com/borongzhang/magpie_ptychography.
@@
  $\bm{z}^{+}_{k}\leftarrow\bm{z}_{k}^{(j)}+\overline{\bm{Q}}\oslash\left(\bm{u}+|\bm{Q}|^{2}\right)\odot\left(\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}\odot\bm{z}_{k}^{(j)}\right)$.
@@
  Coarse probe: $\bm{Q}_{H}\leftarrow\bm{I}_{h}^{H}\bm{Q}$.
  Set weights for downsampling the object: $\bm{W}_{\bm{z}}\leftarrow|\bm{Q}|^{2}\oslash\left(\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}\right)$.
  Set weights for downsampling the revised exit wave: $\bm{W}_{\mathcal{R}}\leftarrow(\bm{I}_{H}^{h}\bm{Q}_{H})\odot\bm{W}_{\bm{z}}\oslash\bm{Q}$.
  Set weights for coarse-grid regularization: $\bm{W}_{\bm{u}}^{H}\leftarrow|\bm{Q}_{H}|^{2}\oslash\left(\bm{I}_{h}^{H}|\bm{Q}|^{2}\right)$.
@@
  $\bm{z}^{+}_{k}\leftarrow\widetilde{\bm{z}}^{(j)}_{k}+\overline{\bm{Q}}\oslash\left(\bm{u}+|\bm{Q}|^{2}\right)\odot\left(\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}\odot\widetilde{\bm{z}}^{(j)}_{k}\right)$.
@@
2:  for $k$ randomly drawn from $1$ to NN without replacement do
@@ 4.2 Converging properties of MAGPIE and weight selection
We summarize our choices of the weight vectors below. These definitions of $\bm{W}_{\bm{z}}$, $\bm{W}_{\mathcal{R}}$, and $\bm{W}_{\bm{u}}^{H}$ ensure the robust performance of MAGPIE.
$\bm{W}_{\bm{z}}=|\bm{Q}|^{2}\oslash\left(\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}\right)\,,\quad\bm{W}_{\mathcal{R}}=(\bm{I}_{H}^{h}\bm{Q}_{H})\odot\bm{W}_{\bm{z}}\oslash\bm{Q}\,,\quad\bm{W}_{\bm{u}}^{H}=|\bm{Q}_{H}|^{2}\oslash\left(\bm{I}_{h}^{H}|\bm{Q}|^{2}\right)\,.$ (16)
As will be shown more clearly in Propositions 4.1, 4.2, and 4.4, and Theorem 4.3. These specific forms not only prevent singularities, maintain coarse-to-fine consistency, and preserve the descent property, but also ensure stability.
It is worth noting that the quantity $\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}$ appears in both $\bm{W}_{\bm{z}}$ and $\bm{W}_{\mathcal{R}}$. Its physical meaning is straightforward: in each $2\times 2$ bin of $|\bm{Q}|^{2}$, $\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}$ computes the average of the four entries and upsamples that average back to the original four positions, whereas $|\bm{Q}|^{2}$ each entry’s original value. As a result,
$\bm{z}_{H,k}^{(j)}=\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\odot\bm{z}_{k}^{(j)}\right)$
@@
To prove this proposition, we realize that $\mathbf{W}_{\mathbf{z}}$ represents the weighted average within a bin and the bin size is $4$. The control for the other two terms comes from Jensen’s inequality used on quadratic functions. The proof is left to Appendix F.
@@
Given the weight vectors defined in  Eqn. (16), the coarse-level objective $\widetilde{\Phi}_{H,k}$ (as defined in  Eqn. (14)) is consistent with the fine-level objective $\widetilde{\Phi}_{k}$ (as defined in  Eqn. (13)) in the sense that, for all $\bm{z}_{k}\in\mathbb{C}^{m^{2}}$, given $\bm{z}_{H,k}=\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\odot\bm{z}_{k}\right)$, we have:
  • $\widetilde{\Phi}_{H,k}\left(\bm{z}_{H,k};\bm{z}_{k}^{(j)}\right)\leq\frac{1}{4}\left\|\bm{W}_{\mathcal{R}}\right\|_{\infty}^{2}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)$ .
  • $\left\|\nabla_{\bm{z}_{H,k}}\widetilde{\Phi}_{H,k}\left(\bm{z}_{H,k};\bm{z}_{k}^{(j)}\right)\right\|_{2}\leq\frac{1}{2}\left\|\bm{W}_{\bm{u}}^{H}\right\|_{\infty}\,\left\|\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)\right\|_{2}$ .
We leave the proof to Appendix G. Most of the proof is a straightforward derivation.
Furthermore, the parameters are chosen so that the coarse-grid solution of Eqn. (14), obtained via the proximal point method, provides a descent direction for the fine-level problem. This is formalized in the following proposition.
@@
Given the weight vectors defined in  Eqn. (16), a proximal point update on the coarse surrogate (14), initialized at $\bm{z}^{\prime}_{H,k}=\bm{I}_{h}^{H}(\bm{W}_{\bm{z}}\odot\bm{z}^{\prime}_{k})$, provides a descent direction for the fine-grid surrogate (13) at $\bm{z}^{\prime}_{k}$. That is:
$\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)^{*}\left(\bm{I}_{H}^{h}\bm{e}_{H,k}\right)\leq 0\,$,
where $\bm{I}_{H}^{h}\bm{e}_{H,k}=\bm{I}_{H}^{h}\left(\widehat{\bm{z}_{H,k}}-\bm{z}_{H,k}^{\prime}\right)$, and $\widehat{\bm{z}_{H,k}}$ solves:
$\widehat{\bm{z}_{H,k}}=\operatornamewithlimits{argmin}_{\bm{z}_{H,k}}\widetilde{\Phi}_{H,k}\left(\bm{z}_{H,k};\bm{z}_{k}^{(j)}\right)+\frac{1}{2}\bm{u}_{H}^{\top}\left|\bm{z}_{H,k}-\bm{z}_{H,k}^{\prime}\right|^{2}\,.$ (17)